HETAV PATEL

Grade 11,

a bar magnet of geometric length 18cm has pole strength 100 A.m. Find magnetic dipole moment of bar magnet

a bar magnet of geometric length 18cm has pole strength 100 A.m. Find magnetic dipole moment of bar magnet
 

Grade:11

1 Answers

Pradeep Tangirala
6 years ago
Magnetic moment = pole strength × 2l
Here 2l means magnetic length...
Now...2l = (5 × L)÷6 
 where L means geometric length... 
On solving...2l = 15 × 10*(-2) = 0.15
So ....M = 100 × 0.15 
                 =15 ampere (meter)*2